How does the strength of the magnetic field that is 1 cm from a current-carrying 

       wire compare with the strength of the magnetic field that is 4 cm from the wire?



1
Expert's answer
2022-01-24T01:13:54-0500

Explanation & Calculations


  • The formula for the magnetic field strength at a point "\small r" far from a straight conductor is

"\qquad\qquad
\begin{aligned}
\small B&=\small \frac{\mu i}{2\pi r}
\end{aligned}"

  • Since this question deals with the same wire with the same current in the same medium ("\small \mu" is constant), this formula can be reduced down to proportional mode as follows,

"\qquad\qquad
\begin{aligned}
\small B&\propto \frac{1}{r} \end{aligned}"

  • Then by comparison, you get,

"\qquad\qquad
\begin{aligned}
\small B_1&\propto \small \frac{1}{1\,cm}\cdots(1)\\
\\
\small B_4&\propto \small \frac{1}{4\,cm}\cdots(2)\\
\\
(2)\,\div\,(1),\\
\\
\small \frac{B_4}{B_1}&=\small \frac{1}{4}\times\frac{1}{1}\\
\small B_4&=\small \frac{1}{4}.B_1
\end{aligned}"

  • Therefore, the field at the far location is one-fourth of that at the near location.